# Convergence to equilibrium for a degenerate three species reaction-diffusion system

### v1

@@ 2. The case of $d_{b}=0$
$\partial_{t}\phi_{\tau}c-d_{c}\Delta\phi_{\tau}c=&\phi^{\prime}_{\tau}c+\phi_{\tau}(ab-c)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \nabla\phi_{\tau}c.\gamma=&0\qquad\qquad\ \ \qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}c(\tau,x)=&0\qquad\qquad\qquad\ \qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
Then from integrability estimation(appendix-11) we have:
$\|\phi_{\tau}c\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)\|\phi^{\prime}_{\tau}c+\phi_{\tau}ab\|_{L^{p}(\Omega_{\tau,\tau+1})}$
(32) $\|c\|_{L^{\infty}(\Omega_{\tau+1,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)(1+M_{\phi})(K_{c}+K_{3})(1+\tau)^{\mu_{c}+\mu_{3}}$
Now again from the relation $\partial_{t}b\leq c$ using jensen inequality we get for some $\mu_{4}\in\mathbb{N}$
$\|b\|_{L^{\infty}(\Omega_{t})}\leq K_{4}(1+t)^{\mu_{4}}$
This along with supremum estimates of $a$((31)) and $c$((32)) gives us the required result, that there exists constant $K_{\infty}$ and $\mu\in\mathbb{N}$, such that:
[…]
Note for $N\geq 4$, if $d_{a},d_{c}$ satisfying closeness condition for $p>N\geq\frac{N+2}{2}$, the assumption in the theorem-4 is automatically satisfied(proposition-1), however the assumption holds true also for $N=1,2,3$ without any closeness condition. ∎
Next we show the integral estimate of $b$, as in lemma-2 and proposition-3, helps us to relate the missing term $\delta_{B}$ with the entropy dissipation functional. The relation is described in the following proposition:
@@
$\eta(t):=\left(\frac{2\min\{d_{a},d_{c},2\}}{P(\Omega)\left(K_{3}+\left|\Omega\right|^{\frac{2}{N}}\right)+1}\right)(1+t)^{-\frac{N-2}{N-1}}$
@@
It can be shown (see [DF06, Lemma 2.1, p.162] for details) that the above defined function satisfies the following bound:
(47) $\Gamma(x,y)\leq C_{\Phi}\max\left\{1,\ln\left(\frac{x}{y}\right)\right\}$
@@
Here the constants $K_{\infty}$ and $\mu$ are the ones appearing in the $\mathrm{L}^{\infty}(\Omega_{t})$ bounds on the concentrations from Lemma 4. The factor $\left(1+\ln(1+t)\right)$ in the above estimate of the relative entropy is due to the fact that Lemma 4 says that at least one of the concentrations has a polynomial (in time) bound on its $\mathrm{L}^{\infty}$-norm. It should be noted that having an uniform (in time) bound on all of the concentrations gets rid of this time factor. In [DF06], the authors prove the following bound
@@
exploiting only the conservation properties (4)-(5). Furthermore, the constant $C_{2}$ in the above bound depends only on the equilibrium states $A_{\infty},B_{\infty},C_{\infty}$ and the constants $M_{1},M_{2}$ from the conservation properties (4)-(5) (see [DF06, Lemma 3.2, p.168] for precise expression for the constant). We thus arrive at the following bound for the relative entropy using (48) and (49):
@@
$\mathcal{S}_{2}$ $=\frac{N-1}{C_{4}(1-\varepsilon)}$
@@
The proof of the above functional inequality is available in [DF06, Lemma 3.3, p.173] which exploits the conservation properties (4) and (5). ∎
@@ 3. The case of $d_{c}=0$
In this section, we shall devote our attention to the study of the degenerate model (3) which corresponds to the vanishing of the diffusion coefficient $d_{c}$. The existence of a smooth positive solution to (3) was proved in [DF15, Theorem 3.2] for all smooth initial non-negative data if the dimension $N\leq 3$. Our main objective of this section is to understand the large time behaviour of solutions to (3). Here, we choose to present our computations when the dimension $N=3$. Analogous results hold true in dimensions one and two as well. Note that we have the entropy equality
@@
It is apparent from the above inequality that the term involving $\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ is missing from its lower bound. A similar scenario was handled in the previous section while dealing with the missing $\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ term. Our strategy was to derive polynomial (in time) bounds on the supremum norms of the concentrations. Here too, we will adapt a similar approach. Note, however, that the proofs of most results in this section markedly differ from the proofs in the previous section. We begin with a time-dependent $\mathrm{L}^{1}(\Omega)$ estimate on certain combinations of the concentrations. This result is inspired by [DFPV07, Theorem 3.1, p.495]. Our proof argues along similar lines as in [DFPV07] while keeping track of the polynomial (in time) bound.
@@
(54) $-\partial_{t}w-Z\Delta w&=H\,\sqrt{Z}\qquad\mbox{ in }\Omega_{T},\\ \nabla w\cdot n(x)&=0\qquad\mbox{ on }\partial\Omega_{T},\\ w(T,x)&=0\qquad\mbox{ for }x\in\Omega.\end{aligned}\right.$
@@
$\int_{0}^{T}\int_{\Omega}\left(a+c\right)^{2}Z\,{\rm d}x\,{\rm d}t\leq\left(2P(\Omega)+8d_{a}T\right)\left\|a_{0}+c_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$
@@
(60) $\left\|a(t,\cdot)\right\|_{\mathrm{L}^{6}(\Omega)}\leq\ell_{6}\left\|a(t,\cdot)\right\|_{\mathrm{H}^{1}(\Omega)},\qquad\left\|b(t,\cdot)\right\|_{\mathrm{L}^{6}(\Omega)}\leq\ell_{6}\left\|b(t,\cdot)\right\|_{\mathrm{H}^{1}(\Omega)}$
@@
$\left\|c(t,\cdot)\right\|^{3}_{\mathrm{L}^{3}(\Omega)}\leq 2^{2}(1+t)^{2}\left(\left\|c_{0}\right\|^{3}_{\mathrm{L}^{3}(\Omega)}+\int_{0}^{t}\left\|ab(s,\cdot)\right\|^{3}_{\mathrm{L}^{3}(\Omega)}\,{\rm d}s\right)$
@@
$\partial_{t}\tilde{a}-d_{a}\Delta\tilde{a}&=0\qquad\mbox{ for }t>0,\,x\in\Omega,\\ \nabla\tilde{a}\cdot n(x)&=0\qquad\mbox{ for }t>0,\,x\in\partial\Omega,\\ \tilde{a}(0,x)&=a_{0}\qquad\mbox{ for }x\in\Omega.\end{aligned}\right.$
We recall the following Gaussian bound on the Neumann Green’s function (see [Mor83, Theorem 2.2, p.37]): there exist positive constants $\texttt{C}_{H},\kappa$ such that
@@
Also see [CK15, Theorem 3.1, p.639] for general parabolic operators. As a consequence we have the following bound on the solution $\tilde{a}$ to the above homogeneous problem:
@@
The proof of the above functional inequality is available in [DF06, Lemma 3.3, p.173] which exploits the conservation properties (4) and (5). ∎
@@
$\partial_{t}\psi+d\Delta\psi&={\color[rgb]{1,0,0}-\theta}\qquad\mbox{ for }(t,x)\in\Omega_{\tau,T},\\ \nabla\psi\cdot n(x)&=0\qquad\mbox{ for }(t,x)\in[\tau,T]\times\partial\Omega,\\ \psi(T,x)&=0\qquad\mbox{ for }x\in\Omega.\end{aligned}\right.$
@@
The proof of (70) can be found in [Lam87, Theorem 1]. We refer to the constant $C_{SOR}$ as the second order regularity constant. Proof of the estimates (71) can be found in [CDF14, Lemma 3.3] and the estimate (72) was derived in [Tan18, Lemma 4.6]. We refer to the constant $C_{IE}$ as the integrability estimation constant.
@@
Proof of the above theorem can be found in [CDF14, Proposition 1.1].

### v2

@@ 2. The case of $d_{b}=0$
$\partial_{t}\phi_{\tau}c-d_{c}\Delta\phi_{\tau}c=&\phi^{\prime}_{\tau}c+\phi_{\tau}(ab-c)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \nabla\phi_{\tau}c.n=&0\qquad\qquad\ \ \qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}c(\tau,x)=&0\qquad\qquad\qquad\ \qquad\ \ \mbox{ in }\Omega\end{aligned}\right.$
Integrability estimation(see Theorem 11 from Appendix) on the above equation yields
$\|\phi_{\tau}c\|_{\mathrm{L}^{\infty}(\Omega_{\tau,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)\|\phi^{\prime}_{\tau}c+\phi_{\tau}ab\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}$
(38) $\|c\|_{\mathrm{L}^{\infty}(\Omega_{\tau+1,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)(1+M_{\phi})(K_{c}+K_{3})(1+\tau)^{2+\mu_{c}}.$
The pointwise estimate of the species $b$ in unit time is a consequence of the above estimate. The relation (36) and the non-negativity of the species $b$, yields
$\|b\|_{\mathrm{L}^{\infty}(\Omega_{\tau})}\leq$ $(1+\tau)\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+\|c\|_{\mathrm{L}^{\infty}(\Omega_{\tau,\tau+1})}(1+\tau)$
$\leq$ $(1+\tau)\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+\bigg{(}C_{IE}(\Omega,N,p)+K_{\infty,[0,2]}\bigg{)}(1+M_{\phi})(K_{c}+K_{3})(1+\tau)^{3+\mu_{c}}.$
By choosing the constant
$K_{\infty}:=\max\left\{\begin{aligned} &K_{\infty,[0,2]},C_{IE}(\Omega,N,p)(1+M_{\phi})K_{c},C_{IE}(\Omega,N,p)(1+M_{\phi})(K_{c}+K_{3}),\\ &\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+\bigg{(}C_{IE}(\Omega,N,p)+K_{\infty,[0,2]}\bigg{)}(1+M_{\phi})(K_{c}+K_{3})\end{aligned}\right\}$
and $\mu=3+\mu_{c}$, we arrive at our result. ∎
Note for the dimension $N\geq 4$, if $d_{a},d_{c}$ satisfying closeness condition (10) for $p>N\geq\frac{N+2}{2}$, the assumption in the Theorem 4 is automatically satisfied (Proposition-1), however the assumption also holds true for $N=1,2,3$ (Proposition 3 and Lemma 3) regardless of any smallness condition on the non-zero diffusion coefficients.
Next we show the integral estimate of $b$, as in Lemma 2 and Proposition-3, helps us to relate the missing term $\delta_{B}$ with the entropy dissipation functional. The relation is described in the following proposition:
@@
$\eta(t):=\left(\frac{2\min\{d_{a},d_{c},2\}}{P(\Omega)\left(K_{3}+\left|\Omega\right|^{\frac{2}{N}}\right)+1}\right)(1+t)^{-\frac{N-2}{N-1}},$
@@
It can be shown (see [4, Lemma 2.1, p.162] for details) that the above defined function satisfies the following bound:
(53) $\Phi(x,y)\leq C_{\Phi}\max\left\{1,\ln\left(\frac{x}{y}\right)\right\}$
@@
Here the constants $K_{\infty}$ and $\mu$ are the ones appearing in the $\mathrm{L}^{\infty}(\Omega_{t})$ bounds on the concentrations from Lemma 4. The factor $\left(1+\ln(1+t)\right)$ in the above estimate of the relative entropy is due to the fact that Lemma 4 says that at least one of the concentrations has a polynomial (in time) bound on its $\mathrm{L}^{\infty}$-norm. It should be noted that having an uniform (in time) bound on all of the concentrations gets rid of this time factor. In [4], the authors prove the following bound
@@
exploiting only the conservation properties (4)-(5). Furthermore, the constant $C_{2}$ in the above bound depends only on the equilibrium states $A_{\infty},B_{\infty},C_{\infty}$ and the constants $M_{1},M_{2}$ from the conservation properties (4)-(5) (see [4, Lemma 3.2, p.168] for precise expression for the constant). We thus arrive at the following bound for the relative entropy using (54) and (55):
@@
$\mathcal{S}_{2}$ $=\frac{N-1}{C_{4}(1-\varepsilon)}.$
@@
The proof of the above functional inequality is available in [4, Lemma 3.3, p.173] which exploits the conservation properties (4) and (5). ∎
@@ 3. The case of $d_{c}=0$
In this section, we shall devote our attention to the study of the degenerate model (3) which corresponds to the vanishing of the diffusion coefficient $d_{c}$. The existence of a smooth positive solution to (3) was proved in [5, Theorem 3.2] for all smooth initial non-negative data if the dimension $N\leq 3$. Our main objective of this section is to understand the large time behaviour of solutions to (3). Here, we choose to present our computations when the dimension $N=3$. Analogous results hold true in dimensions one and two as well. Note that we have the entropy equality
@@
It is apparent from the above inequality that the term involving $\left\|\delta_{C}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ is missing from its lower bound. A similar scenario was handled in the previous section while dealing with the missing $\left\|\delta_{B}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}$ term. Our strategy was to derive polynomial (in time) bounds on the supremum norms of the concentrations. Here too, we will adapt a similar approach. Note, however, that the proofs of most results in this section markedly differ from the proofs in the previous section. We begin with a time-dependent $\mathrm{L}^{1}(\Omega)$ estimate on certain combinations of the concentrations. This result is inspired by [6, Theorem 3.1, p.495]. Our proof argues along similar lines as in [6] while keeping track of the polynomial (in time) bound.
@@
(60) $-\partial_{t}w-Z\Delta w&=H\,\sqrt{Z}\qquad\mbox{ in }\ \Omega_{T},\\ \nabla w\cdot n(x)&=0\qquad\qquad\mbox{ on }\ \partial\Omega_{T},\\ w(T,x)&=0\qquad\qquad\ \mbox{ in }\ \Omega.\end{aligned}\right.$
@@
$\int_{0}^{T}\int_{\Omega}\left(a+c\right)^{2}Z\,{\rm d}x\,{\rm d}t\leq\left(2P(\Omega)+8d_{a}T\right)\left\|a_{0}+c_{0}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$
@@
(66) $\left\|a(t,\cdot)\right\|_{\mathrm{L}^{6}(\Omega)}\leq\ell_{6}\left\|a(t,\cdot)\right\|_{\mathrm{H}^{1}(\Omega)},\qquad\left\|b(t,\cdot)\right\|_{\mathrm{L}^{6}(\Omega)}\leq\ell_{6}\left\|b(t,\cdot)\right\|_{\mathrm{H}^{1}(\Omega)}.$
@@
$\left\|c(t,\cdot)\right\|^{3}_{\mathrm{L}^{3}(\Omega)}\leq 2^{2}(1+t)^{2}\left(\left\|c_{0}\right\|^{3}_{\mathrm{L}^{3}(\Omega)}+\int_{0}^{t}\left\|ab(s,\cdot)\right\|^{3}_{\mathrm{L}^{3}(\Omega)}\,{\rm d}s\right).$
@@
$\partial_{t}\tilde{a}-d_{a}\Delta\tilde{a}&=0\qquad\ \mbox{ in }\ \Omega_{t},\\ \nabla\tilde{a}\cdot n(x)&=0\qquad\ \mbox{ on }\ \partial\Omega_{t},\\ \tilde{a}(0,x)&=a_{0}\qquad\mbox{ in }\ \Omega.\end{aligned}\right.$
We recall the following Gaussian bound on the Neumann Green’s function (see [10, Theorem 2.2, p.37]): there exist positive constants $\texttt{C}_{H},\kappa$ such that
@@
Also see [3, Theorem 3.1, p.639] for general parabolic operators. As a consequence we have the following bound on the solution $\tilde{a}$ to the above homogeneous problem:
@@
The proof of the above functional inequality is available in [4, Lemma 3.3, p.173] which exploits the conservation properties (4) and (5). ∎
@@
$\partial_{t}\psi+d\Delta\psi&=-\theta\qquad\mbox{ for }(t,x)\in\Omega_{\tau,T},\\ \nabla\psi\cdot n(x)&=0\qquad\ \ \mbox{ for }(t,x)\in[\tau,T]\times\partial\Omega,\\ \psi(T,x)&=0\qquad\ \ \mbox{ for }x\in\Omega.\end{aligned}\right.$
@@
The proof of (76) can be found in [9, Theorem 1]. We refer to the constant $C_{SOR}$ as the second order regularity constant. Proof of the estimates (77) can be found in [2, Lemma 3.3] and the estimate (78) was derived in [15, Lemma 4.6]. We refer to the constant $C_{IE}$ as the integrability estimation constant.
@@
Proof of the above theorem can be found in [2, Proposition 1.1].